OpenStudy (anonymous):

A paint can is in the shape of a right circular cylinder. The volume of the paint can is 600 cubic inches and its altitude is 12 inches. Find the radius, in inches, of the base of the paint can. Express the answer in simplest radical form. Im studying for a regents i have a answer but i need it explained

OpenStudy (anonymous):

answer is \[5\sqrt{2} \] and 533.1 as stated by my answer sheet.

jhonyy9 (jhonyy9):

so do you know what is the formula for volum of a cylinder ?

OpenStudy (anonymous):

v=BH where B is the area of the base

jhonyy9 (jhonyy9):

yes right and area of base from what terms can you calcule it ? what is this formula for area of base ?

OpenStudy (anonymous):

area of the base would be pi r squared?

jhonyy9 (jhonyy9):

yes right so than how can you writing the volum of cylinder using formula for area of base ?

jhonyy9 (jhonyy9):

V=?

OpenStudy (anonymous):

could i do 600pi = BH(12) ? then find base? the volume was given as 600pi

jhonyy9 (jhonyy9):

what is there (12) ?

OpenStudy (anonymous):

12 is altitude

OpenStudy (anonymous):

they want me to find radius so i can find the area

jhonyy9 (jhonyy9):

so yes is right but when you write formula for volum you need writing just fonts or numbers - so volum of cylinder is pi*r squared time height so V=pir^2 *h -yes is right - so from this you can calcule r yes ?

OpenStudy (anonymous):

checking

OpenStudy (anonymous):

i dont understand this question

jhonyy9 (jhonyy9):

so you need calcule radius of circle of the base of cylinder ,yes ?

jhonyy9 (jhonyy9):

so ok V=pi*r^2 *h 600pi=pi*r^2 *12 so divide both sides by 12pi and will get 600pi 12pi*r^2 ------ = ------- 12pi 12pi 50=r^2 so because a length not can being negativ value so than r=sqrt50 =5sqrt2 r=5sqrt2

jhonyy9 (jhonyy9):

hope so much that is understandably sure right for you good luck bye